There is one and only one complex cube root of 1.
True or false with correct explanation
1
Expert's answer
2020-10-12T05:41:38-0400

False.


There are three complex cube roots of 1.


"1>0" so we can find these cube roots:

"x_1=\sqrt{1}=1"

"x_2=\sqrt{1}*(-\dfrac{1}{2}+\dfrac{\sqrt{3}}{2}i)=-\dfrac{1}{2}+\dfrac{\sqrt{3}}{2}i"

"x_3=\sqrt{1}*(-\dfrac{1}{2}-\dfrac{\sqrt{3}}{2}i)=-\dfrac{1}{2}-\dfrac{\sqrt{3}}{2}i"

where

"\sqrt{1}=1" is a simple cube root of 1.


"(\sqrt{1})^3=1"

"(-\dfrac{1}{2}+\dfrac{\sqrt{3}}{2}i)^2=\dfrac{\sqrt{3}}{2}i-\dfrac{2}{4}"

"(-\dfrac{1}{2}+\dfrac{\sqrt{3}}{2}i)*(\dfrac{\sqrt{3}}{2}i-\dfrac{2}{4})=1"